OpenStudy (anonymous):

Rewrite with only sin x and cos x. sin 3x

jhonyy9 (jhonyy9):

to the first step us sin(a+b)= ? for sin(2x+x) = ?

OpenStudy (anonymous):

yes i realize that, then i think it is supposed to be sinxcos2x + cosxsin2x right?

jhonyy9 (jhonyy9):

yes and for this you know again formula for sin2x = ? and for cos2x = ?

OpenStudy (anonymous):

ok so cosx(2sinxcosx) + sinx(2cos^2(x) - 1 or is there a better double angle formula to use for the cos2x?

jhonyy9 (jhonyy9):

not is necessary so after this you know again formula sni^2 x +cos^2 x = 1 us this and finish

OpenStudy (anonymous):

ok now im confused, where do you use that formula

jhonyy9 (jhonyy9):

ok so what you have got above continue it eleminnate the parantheses

jhonyy9 (jhonyy9):

see than what you will get

OpenStudy (anonymous):

2sinxcos^2x + 2cos^2xsinx-sinx

OpenStudy (anonymous):

is that right so far?

jhonyy9 (jhonyy9):

how you have got -sinx from the end ?

OpenStudy (anonymous):

i multiplied the -1 by the sinx when i distributed it

jhonyy9 (jhonyy9):

yes ok but there from the end you have missed one paranthese

OpenStudy (anonymous):

I don't understand, what should i be at now?

jhonyy9 (jhonyy9):

so you have got this 2sinxcos^2x + 2cos^2xsinx-sinx ok and now what you need doing again ?

OpenStudy (anonymous):

thats where i get to every time and dont know what to do afterwards

jhonyy9 (jhonyy9):

so check it above what saying the text of this exercise

jhonyy9 (jhonyy9):

do you understand it now ?

OpenStudy (anonymous):

not really, do i need to simplify the 2sinx?

jhonyy9 (jhonyy9):

if you read the text of this exercise so thgis saying that you need rewriting the sin3x using sinx and cosx yes ?

OpenStudy (anonymous):

yes

jhonyy9 (jhonyy9):

so if you check on the final line what you have got ?

OpenStudy (anonymous):

i know that, but it still isnt one of my answers

jhonyy9 (jhonyy9):

2sinxcos^2x + 2cos^2xsinx-sinx what you can doing again here ?

OpenStudy (anonymous):

i don't know, i don't see where i can simplify anything else

jhonyy9 (jhonyy9):

why ? the first and second terms not are the same ?

jhonyy9 (jhonyy9):

2sinxcos^2x not is equale by 2cos^2xsinx ?

OpenStudy (anonymous):

i know so multiply the right side by 2?

jhonyy9 (jhonyy9):

2sinxcos^2x + 2cos^2xsinx-sinx let sinx=a and cosx=b than there is 2ab^2 +2b^2a -a = ?

OpenStudy (anonymous):

yes

jhonyy9 (jhonyy9):

so how do you think ab^2 not is equal b^2a ?

OpenStudy (anonymous):

it is i guess i was seeing it the wrong way. but now i do see how they are equal

jhonyy9 (jhonyy9):

so than now you have got the right answer ?

OpenStudy (anonymous):

no, it still isn't any of the answers given to me

OpenStudy (anonymous):

2 sin x cos2x + cos x 2 sin x cos2x + sin3x sin x cos2x - sin3x + cos3x 2 cos2x sin x + sin x - 2 sin3x

OpenStudy (anonymous):

these are my choices

jhonyy9 (jhonyy9):

2sinxcos^2x + 2cos^2xsinx-sinx = 2sinxcos^2x + 2sinxcos^2x -sinx = =

OpenStudy (anonymous):

uhm...4sinxcox^2x-sinx

jhonyy9 (jhonyy9):

yes and continue it

jhonyy9 (jhonyy9):

factoriz out the sinx

OpenStudy (anonymous):

sinx(3sinxcos^2x-1)

jhonyy9 (jhonyy9):

how you have got 3sinxcos^2x

OpenStudy (anonymous):

cuz i took the sinx out

jhonyy9 (jhonyy9):

2sinxcos^2x + 2cos^2xsinx-sinx = =4sinxcos^2x -sinx = =sinx(4cos^2x -1) = = ?

OpenStudy (anonymous):

oh ok now i see what you did

jhonyy9 (jhonyy9):

ok but for 4cos^2x -1 = ? us formula a^2 -b^2 = ?

OpenStudy (anonymous):

i dont know

jhonyy9 (jhonyy9):

a^2 -b^2 = (a-b)(a+b)

OpenStudy (anonymous):

ok so (2cosx-1)(2cosx+1)

jhonyy9 (jhonyy9):

2sinxcos^2x + 2cos^2xsinx-sinx = =4sinxcos^2x -sinx = =sinx(4cos^2x -1) = =sinx(4cos^2x -(sin^2x +cos^2x) )= ?

OpenStudy (anonymous):

do we then simplify the sin^2x + cox^2x

jhonyy9 (jhonyy9):

=sinx(4cos^x -sin^2x -cos^2x) =

jhonyy9 (jhonyy9):

sorry there is 4cos^2x

OpenStudy (anonymous):

then do you distribute the sinx?

jhonyy9 (jhonyy9):

yes but before you can assuming 4cos^2x -cos^2x = ?

OpenStudy (anonymous):

sinx(3cos^2x-sin^2x)

jhonyy9 (jhonyy9):

yes and continue

OpenStudy (anonymous):

3cos^2xsinx-sin^3x

jhonyy9 (jhonyy9):

yes

OpenStudy (anonymous):

but that isn't one of my answers